OpenStudy (anonymous):

The sum of the first hundred terms of an arithmetic progression with first term 'a' and common difference 'd' is T. The sum of the first 50 odd-numbered terms, i.e. the first, third, fifth,........, ninety-ninth is [(T/2) - 100]. Find the value of a.

hartnn (hartnn):

it should be 100a in equation 1 ,isn't it?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

S(total)=100/2(2a+99d)=T = 100a+4950d=T ------------- (1) 1,3,5,7,........, note:common diff.=2d S(odd) = 50/2(2a+98d)=T/2-100 simplify it = 100a+9800d+200=T-------------(2) Solve (1) and (2)

OpenStudy (anonymous):

Then I got: d= - (4/97)

OpenStudy (anonymous):

is it correct ans.?

OpenStudy (anonymous):

substitute d in (1)

OpenStudy (anonymous):

but there is again two unknown variable.. i.e. T and a.... and again it the result is an equation.

OpenStudy (anonymous):

so put d in both and solve both the resulting eq.

hartnn (hartnn):

u want 3rd equation right?so that u can have 3 equation ,3 unknowns to get a value? here's how u get 3rd eq: since, sum of odd numbered term + sum of even numbered term = total sum so u get even numbered term as T-(T/2-100)=T/2 + 100 with this u can form 3rd equation as 50/28(2(a+d)+98d)=T/2+100 because,1st term here is a+d now can u solve these 3 equations with 3 unknowns to get a ??

hartnn (hartnn):

@curiousshubham

hartnn (hartnn):

* 50/2*(2(a+d)+98d)=T/2+100

OpenStudy (anonymous):

Just I checked the answer key and I found the answer as 40 i.e. a = 40.

hartnn (hartnn):

in the 2nd equation above there should be 4900d instead of 9800d

hartnn (hartnn):

to get d=4

OpenStudy (anonymous):

@hartnn I am really confused.......

hartnn (hartnn):

@curiousshubham are u sure about the question??is it to find a or to find d only? and d we easily found out = 4 but we need an extra equation/information to find a,T

OpenStudy (anonymous):

@hartnn I am sure about the question and its in my coursebook. I checked the question several times to be sure.

hartnn (hartnn):

then we both need to seek help from a smarter mathematician :P maybe @mukushla can help...

OpenStudy (anonymous):

@hartnn no problem with ur work...